The total area of a regular tetrahedron is 110.85 m2, determine its base edge.

6. If the lateral area of a right circular cylinder is 90 cm2 and its volume is 225 cm3, find its radius.

7. Determine the lateral area of a cylinder of height 300mm and base radius of 100mm.
1
Expert's answer
2020-12-15T02:16:44-0500

5).

A = "\sqrt3" a2

a2 = "\frac{A}{\sqrt3}"


a2 = "\frac{110.85}{\sqrt3}"


a = 7.99995m

a = 8m


6).

Area of right circular cylinder is "A= 2\pi r h" and volume is "V= \pi r^2 h" . Thus we have

"2\pi rh= 90 \\
rh =\cfrac{45}{\pi}......Eq[1]\\ 
r^2 h=\cfrac{225}{\pi}.......Eq[2]"

Now dividing Eq[2] by Eq[1]...

"r=\cfrac{225}{45}\\
r= 5cm.....Ans"

radius of right circular cylinder is 5 cm.


7).

"h = 300\; mm \\

r = 100 \;mm \\

Lateral \;area\; of \;cylinder = 2πrh \\

= 2 \times 3.14 \times 100 \times 300 \\

= 188520 \; mm^2"

Answer: 188520 mm2